# KiDS-Legacy: Consistency of cosmic shear measurements and joint cosmological constraints with external probes

### v1

@@ Appendix A Catalogue level splits: data properties and B-mode analysis
For the colour-based split, we define red galaxies via a selection on the spectral type $T_{\rm B}$ reported by the bpz code. We adopt a threshold of $T_{\rm B}\leq 1.9$ from appendix B of W25, which selects objects with contributions from an elliptical galaxy spectrum (template E1). For this split, the red galaxy sample only encompasses the first five tomographic bins since the catalogue only contains very few red galaxies with $z_{\rm B}>1.14$, making it challenging to calibrate the redshift distribution of the sixth bin. However, given the vanishingly low signal-to-noise ratio of such a sparsely populated bin, we do not expect the exclusion of this bin to make an impact on the analysis. By contrast, Li et al. (2021) previously performed a consistency test between red and blue galaxies with cosmic shear data from the third KiDS data release (KV450, de Jong et al. 2017; Wright et al. 2019), defining blue galaxies via a threshold of $T_{\rm B}\leq 3.0$. This selection additionally encompasses objects with contributions from two spiral galaxy templates (Sbc and Sdc), while objects with contributions of irregular and starburst galaxy spectra (templates Im, SB2, and SB3) are labelled as blue galaxies. Originally, this threshold was chosen in order to ensure a similar constraining power per data subset, although this no longer holds in KiDS-Legacy given the addition of high redshift galaxies, which are predominantly blue.
Figure 12: Redshift distributions per tomographic bin for catalogue level splits. Left panel: North-South split. Middle panel: Red-blue split defined via a cut on the spectral type of $T_{\rm B}=3.0$. Right panel: Red-blue split defined via a cut on the spectral type of $T_{\rm B}=1.9$.
Table 6: Data properties per tomographic bin for catalogue level splits.
@@
5 $0.90<z_{\rm B}\leq 1.14$ 15.8 1.35 0.28 $-0.011\pm 0.010$ $\phantom{-}0.030\pm 0.008$
6 $1.14<z_{\rm B}\leq 2.00$ 14.6 1.07 0.30 $-0.054\pm 0.011$ $\phantom{-}0.045\pm 0.009$
North 1 $0.10<z_{\rm B}\leq 0.42$ 09.1 1.74 0.28 $-0.025\pm 0.010$ $-0.024\pm 0.009$
2 $0.42<z_{\rm B}\leq 0.58$ 09.4 1.67 0.27 $\phantom{-}0.013\pm 0.010$ $-0.021\pm 0.009$
3 $0.58<z_{\rm B}\leq 0.71$ 08.6 1.50 0.29 $-0.001\pm 0.010$ $-0.008\pm 0.011$
@@
5 $0.90<z_{\rm B}\leq 1.14$ 07.7 1.27 0.28 $-0.011\pm 0.010$ $\phantom{-}0.021\pm 0.011$
6 $1.14<z_{\rm B}\leq 2.00$ 06.9 0.99 0.30 $-0.057\pm 0.011$ $\phantom{-}0.040\pm 0.013$
South 1 $0.10<z_{\rm B}\leq 0.42$ 09.0 1.81 0.28 $-0.025\pm 0.010$ $-0.017\pm 0.010$
2 $0.42<z_{\rm B}\leq 0.58$ 08.6 1.62 0.27 $\phantom{-}0.013\pm 0.010$ $-0.010\pm 0.010$
3 $0.58<z_{\rm B}\leq 0.71$ 08.1 1.49 0.29 $-0.001\pm 0.010$ $-0.015\pm 0.012$
@@
5 $0.90<z_{\rm B}\leq 1.14$ 08.1 1.43 0.28 $-0.012\pm 0.010$ $\phantom{-}0.026\pm 0.012$
6 $1.14<z_{\rm B}\leq 2.00$ 07.6 1.16 0.30 $-0.057\pm 0.011$ $\phantom{-}0.047\pm 0.013$
Red: $T_{\rm B}\leq 3.0$ 1 $0.10<z_{\rm B}\leq 0.42$ 06.8 0.73 0.27 $-0.008\pm 0.010$ $-0.031\pm 0.011$
2 $0.42<z_{\rm B}\leq 0.58$ 06.0 0.60 0.25 $-0.003\pm 0.010$ $-0.038\pm 0.011$
3 $0.58<z_{\rm B}\leq 0.71$ 06.1 0.62 0.29 $-0.016\pm 0.010$ $-0.043\pm 0.013$
4 $0.71<z_{\rm B}\leq 0.90$ 06.7 0.62 0.25 $-0.003\pm 0.010$ $-0.017\pm 0.011$
5 $0.90<z_{\rm B}\leq 1.14$ 06.4 0.58 0.30 $-0.014\pm 0.010$ $-0.043\pm 0.014$
6 $1.14<z_{\rm B}\leq 2.00$ 02.7 0.21 0.30 $-0.081\pm 0.011$ $-0.032\pm 0.020$
Blue: $T_{\rm B}>3.0$ 1 $0.10<z_{\rm B}\leq 0.42$ 11.3 1.00 0.28 $-0.025\pm 0.010$ $-0.016\pm 0.007$
2 $0.42<z_{\rm B}\leq 0.58$ 12.0 1.03 0.28 $\phantom{-}0.030\pm 0.010$ $-0.005\pm 0.007$
@@
Blue: $T_{\rm B}>1.9$ 1 $0.10<z_{\rm B}\leq 0.42$ 15.5 1.48 0.28 $-0.023\pm 0.010$ $-0.022\pm 0.006$
2 $0.42<z_{\rm B}\leq 0.58$ 14.5 1.29 0.28 $\phantom{-}0.024\pm 0.010$ $-0.010\pm 0.006$
3 $0.58<z_{\rm B}\leq 0.71$ 13.7 1.20 0.29 $\phantom{-}0.007\pm 0.010$ $-0.001\pm 0.008$
@@
5 $0.90<z_{\rm B}\leq 1.14$ 12.8 1.08 0.27 $-0.019\pm 0.010$ $\phantom{-}0.044\pm 0.008$
6 $1.14<z_{\rm B}\leq 2.00$ 14.1 1.03 0.30 $-0.062\pm 0.011$ $\phantom{-}0.048\pm 0.009$
 [^13] We list the index of the tomographic bin, the range in photometric redshift $z_{\rm B}$, the fraction of sources with respect to the total number of sources, the effective number density $n_{\rm eff}$, the ellipticity dispersion $\sigma_{\rm e}$, the shift in the mean of the redshift distribution $\delta_{z}$, and the multiplicative shear bias $m$.
To first order, we expect cosmic shear to only produce E-mode signals, making B-mode signals negligible for current stage-III surveys. Therefore, B modes are a useful quantity for null-tests for residual systematics in the cosmic shear measurement. Here, we quantify the significance of the B modes for the split catalogues and refer to appendix D in W25 for a discussion of B modes in the fiducial catalogue. We employ COSEBIs as test statistic since it allows for a clean separation of E and B modes. To quantify the significance of the B mode, we compute the $\chi^{2}$ assuming the null hypothesis and compute the corresponding $p$-value. Here, the $p$-value is equal to the probability of producing a B mode that is more significant than the observed signal, assuming that the B mode is randomly drawn from a Gaussian distribution with zero mean.
For each subset of the catalogue, we compute the first six COSEBIs B modes and quantify the significance for the full data vector consisting of all 21 combinations of the six tomographic redshift bins. Additionally, we quantify the B mode significance individually for each tomographic bin combination. The measured B-mode signal for the three catalogue level splits are presented in Figs. 13, 14, and 15. All $p$-values pass our required threshold of $p>0.01$, which is the community standard for B-mode tests (Dark Energy Survey and Kilo-Degree Survey Collaboration et al. 2023), for both the individual tomographic bin combination and for the combination of all tomographic bins.
Figure 13: COSEBIs B-mode measurements for the North-South split catalogue. The green and purple data points show the measurements of the KiDS-North and KiDS-South sample, respectively. Each panel represents auto- or cross-correlation between tomographic bins, as indicated by the label in the top right corner. The corresponding $p$-value is denoted in the top left corner of each panel. The $p$-value of the combined data vector is given in the top left corner of the figure. For visualisation purposes, we display the discrete $n$ modes with an offset on the x-axis. We note that the B-mode signals are highly correlated within a tomographic bin and advise against a so-called ?$\chi$-by-eye?.
Figure 14: Same as Fig. 13, but for the red-blue split catalogue defined via a cut on the spectral type $T_{\rm B}=3.0$.
Figure 15: Same as Fig. 13, but for the red-blue split catalogue defined via a cut on the spectral type $T_{\rm B}=1.9$. For this threshold, the catalogue only contains very few red galaxies with $z_{\rm B}>1.14$. Therefore, the red galaxy sample only encompasses the first five tomographic bins.
@@ Appendix B Effective number of constrained parameters
For the calculation of the tension probability with the suspiciousness statistic, we require the difference between the effective number of constrained parameters $N_{\Theta}$ in the fiducial and split cosmological model, as defined in Eq. (7). We follow the methodology of Joachimi et al. (2021) and infer $N_{\Theta}$ from mock realisations of the cosmic shear data vector. Assuming a Planck Collaboration et al. (2020) cosmology and our fiducial data covariance matrix, we generate 1000 realisations of the data vector from a multivariate Gaussian distribution. For each mock realisation, we then maximise the posterior and obtain an estimate of the best-fit $\chi^{2}$. As discussed in Joachimi et al. (2021), the distribution of $\chi^{2}_{\rm best}$ is well described by a $\chi^{2}$-distribution with a number of degrees of freedom $k_{\rm dof}=N_{\rm d}-N_{\Theta,\chi^{2}}$, where $N_{\rm d}$ denotes the dimensionality of the data vector. We therefore fit the distribution of $\chi^{2}_{\rm best}$-values from mock data vectors to a $\chi^{2}$-distribution in order to determine the effective number of degrees of freedom, denoted $N_{\Theta,\chi^{2}}$.
The inferred estimates of $N_{\Theta}$ are summarised in Table 7 for the fiducial cosmic shear analysis setup with all summary statistics, dubbed ?1cosmo?, as well as for the split cosmological analysis setups with COSEBIs considered in Sect. 4. For the split between angular scales, we employ band powers and 2PCFs, as discussed in Sect. 4.1.3. Additionally, we infer $N_{\Theta}$ for catalogue- and statistic-level splits in a single cosmological setup. For comparison, we provide the Bayesian model dimensionality $N_{\Theta,{\rm BMD}}$ computed via Eq. (8). For both estimates, we list the resulting difference in $N_{\Theta}$ between the fiducial cosmological model and the split cosmological model, denoted by $d_{\chi^{2}}$ and $d_{\rm BMD}$. Overall, we find the number of constrained parameters to be higher in the split cosmological model. We attribute this to the doubling of the cosmological parameter space, where both instances of the parameters to which cosmic shear is sensitive are independently constrained.
Table 7: Effective number of constrained parameters
@@
Catalogue North vs south 9.17 7.45 3.03 1.81
@@
Statistic COSEBIs vs Band powers 10.22 10.69 3.26 4.59
@@
 [^14] For each analysis setup we provide the effective number of constrained parameters inferred via $\chi^{2}$ minimisation of mock data vectors (Joachimi et al. 2021) in the third column, and the Bayesian model dimensionality (Handley & Lemos 2019a), inferred directly from the chain, in the fourth column. The first three rows show the results in an analysis with the fiducial setup, dubbed ‘1cosmo‘, in which we do not apply any splits to the data. The remaining rows present results when splitting the data into subsets. Here, we report the effective number of constrained parameters with the split cosmological model. For the splits at the catalogue level and the splits by summary statistic we additionally report the results with a shared set of parameters. The final two columns report the difference in the number of constrained parameters between the fiducial cosmological model and the split cosmological model. We employ COSEBIs as the summary statistic except when explicitly stated.

### v2

@@ Appendix A Catalogue-level splits: Data properties and B-mode analysis
For the colour-based split, we defined red galaxies via a selection on the spectral type $T_{\rm B}$ reported by the bpz code. We adopted a threshold of $T_{\rm B}\leq 1.9$ from appendix B in W25, which selects objects with contributions from an elliptical galaxy spectrum (template E1). For this split, the red galaxy sample only encompasses the first five tomographic bins since the catalogue only contains very few red galaxies with $z_{\rm B}>1.14$, making it challenging to calibrate the redshift distribution of the sixth bin. However, given the vanishingly low signal-to-noise ratio of such a sparsely populated bin, we do not expect the exclusion of this bin to make an impact on the analysis. By contrast, Li et al. (2021) previously performed a consistency test between red and blue galaxies with cosmic shear data from the third KiDS data release (KV450, de Jong et al. 2017; Wright et al. 2019), defining blue galaxies via a threshold of $T_{\rm B}\leq 3.0$. This selection additionally encompasses objects with contributions from two spiral galaxy templates (Sbc and Sdc), while objects with contributions of irregular and starburst galaxy spectra (templates Im, SB2, and SB3) are labelled as blue galaxies. Originally, this threshold was chosen in order to ensure a similar constraining power per data subset, although this no longer holds in KiDS-Legacy given the addition of high redshift galaxies, which are predominantly blue.
Figure 12: Redshift distributions per tomographic bin for catalogue-level splits. Left panel: North-South split. Middle panel: Red-blue split defined via a cut on the spectral type of $T_{\rm B}=3.0$. Right panel: Red-blue split defined via a cut on the spectral type of $T_{\rm B}=1.9$.
Table 6: Data properties per tomographic bin for catalogue-level splits. [^13] We list the index of the tomographic bin, the range in photometric redshift $z_{\rm B}$, the fraction of sources with respect to the total number of sources, the effective number density $n_{\rm eff}$, the ellipticity dispersion $\sigma_{\rm e}$, the shift in the mean of the redshift distribution $\delta_{z}$, and the multiplicative shear bias $m$.
@@
[…]
North 1 $0.10<z_{\rm B}\leq 0.42$ 09.1 1.74 0.28 $025\pm 0.010 0.0$$024\pm 0.009$
2 $0.42<z_{\rm B}\leq 0.58$ 09.4 1.67 0.27 $\phantom{-}0.013\pm 0.010$ $0.021\pm 0.009$
3 $0.58<z_{\rm B}\leq 0.71$ 08.6 1.50 0.29 $001\pm 0.010 0.0$$008\pm 0.011$
@@
5 $0.90<z_{\rm B}\leq 1.14$ 07.7 1.27 0.28 $011\pm 0.010 0.0$$1\phantom{-}0.021\pm$ 0.011
[…]
South 1 $0.10<z_{\rm B}\leq 0.42$ 09.0 1.81 0.28 $025\pm 0.010 0.0$$017\pm 0.010$
2 $0.42<z_{\rm B}\leq 0.58$ 08.6 1.62 0.27 $\phantom{-}0.013\pm 0.010$ $0.010\pm 0.010$
3 $0.58<z_{\rm B}\leq 0.71$ 08.1 1.49 0.29 $001\pm 0.010 0.0$$015\pm 0.012$
@@
[…]
Red: $T_{\rm B}\leq 3.0$ 1 $0.10<z_{\rm B}\leq 0.42$ 06.8 0.73 0.27 $008\pm 0.010 0.0$$031\pm 0.011$
2 $0.42<z_{\rm B}\leq 0.58$ 06.0 0.60 0.25 $003\pm 0.010 0.0$$038\pm 0.011$
3 $0.58<z_{\rm B}\leq 0.71$ 06.1 0.62 0.29 $016\pm 0.010 0.0$$043\pm 0.013$
4 $0.71<z_{\rm B}\leq 0.90$ 06.7 0.62 0.25 $003\pm 0.010 0.0$$017\pm 0.011$
5 $0.90<z_{\rm B}\leq 1.14$ 06.4 0.58 0.30 $014\pm 0.010 0.0$$043\pm 0.014$
6 $1.14<z_{\rm B}\leq 2.00$ 02.7 0.21 0.30 $081\pm 0.011 0.0$$032\pm 0.020$
Blue: $T_{\rm B}>3.0$ 1 $0.10<z_{\rm B}\leq 0.42$ 11.3 1.00 0.28 $025\pm 0.010 0.0$$016\pm 0.007$
2 $0.42<z_{\rm B}\leq 0.58$ 12.0 1.03 0.28 $\phantom{-}0.030\pm 0.010$ $0.005\pm 0.007$
@@
Blue: $T_{\rm B}>1.9$ 1 $0.10<z_{\rm B}\leq 0.42$ 15.5 1.48 0.28 $3\pm 0.010 0.022±0.0$$2\pm 0.006$
2 $0.42<z_{\rm B}\leq 0.58$ 14.5 1.29 0.28 $\phantom{-}0.024\pm 0.010$ $0.010\pm 0.006$
3 $0.58<z_{\rm B}\leq 0.71$ 13.7 1.20 0.29 $\phantom{-}0.007\pm 0.010$ $0.001\pm 0.008$
@@
[…]
Figure 13: COSEBIs B-mode measurements for the North-South split catalogue. The green and purple data points show the measurements of the KiDS-North and KiDS-South sample, respectively. Each panel represents auto- or cross-correlation between tomographic bins, as indicated by the label in the top right corner. The corresponding $p$-value is denoted in the top left corner of each panel. The $p$-value of the combined data vector is given in the top left corner of the figure. For visualisation purposes, we display the discrete $n$ modes with an offset on the x-axis. We note that the B-mode signals are highly correlated within a tomographic bin and advise against a so-called $\chi$-by-eye?.
Figure 14: Same as Fig. 13 but for the red-blue split catalogue defined via a cut on the spectral type $T_{\rm B}=3.0$.
Figure 15: Same as Fig. 13 but for the red-blue split catalogue defined via a cut on the spectral type $T_{\rm B}=1.9$. For this threshold, the catalogue only contains very few red galaxies with $z_{\rm B}>1.14$. Therefore, the red galaxy sample only encompasses the first five tomographic bins.
To the first order, we expect cosmic shear to only produce E-mode signals, making B-mode signals negligible for current stage-III surveys. Therefore, B modes are a useful quantity for null-tests for residual systematics in the cosmic shear measurement. Here, we quantified the significance of the B modes for the split catalogues, referring to appendix E in W25 for a discussion of B modes in the fiducial catalogue. We employed COSEBIs as test statistic since it allows for a clean separation of E and B modes. To quantify the significance of the B mode, we computed the $\chi^{2}$ values assuming the null hypothesis and compute the corresponding $p$-value. Here, the $p$-value is equal to the probability of producing a B mode that is more significant than the observed signal, assuming that the B mode is randomly drawn from a Gaussian distribution with zero mean.
For each subset of the catalogue, we computed the first six COSEBIs B modes and quantified the significance for the full data vector consisting of all 21 combinations of the six tomographic redshift bins. Additionally, we quantified the B mode significance individually for each tomographic bin combination. The measured B-mode signal for the three catalogue-level splits are presented in Figs. 13, 14, and 15. All $p$-values pass our required threshold of $p>0.01$, which is the community standard for B-mode tests (Dark Energy Survey and Kilo-Degree Survey Collaboration et al. 2023), for both the individual tomographic bin combination and for the combination of all tomographic bins.
@@ Appendix B Effective number of constrained parameters
For the calculation of the tension probability with the suspiciousness statistic, we require the difference between the effective number of constrained parameters $N_{\Theta}$ in the fiducial and split cosmological model, as defined in Eq. (7). We followed the methodology of Joachimi et al. (2021) and infer $N_{\Theta}$ from mock realisations of the cosmic shear data vector. Assuming a Planck Collaboration et al. (2020) cosmology and our fiducial data covariance matrix, we generated 1000 realisations of the data vector from a multivariate Gaussian distribution. For each mock realisation, we then maximised the posterior and obtain an estimate of the best-fit $\chi^{2}$. As discussed in Joachimi et al. (2021), the distribution of $\chi^{2}_{\rm best}$ is well described by a $\chi^{2}$-distribution with a number of degrees of freedom $k_{\rm dof}=N_{\rm d}-N_{\Theta,\chi^{2}}$, where $N_{\rm d}$ denotes the dimensionality of the data vector. We therefore fit the distribution of $\chi^{2}_{\rm best}$-values from mock data vectors to a $\chi^{2}$-distribution to determine the effective number of degrees of freedom, denoted $N_{\Theta,\chi^{2}}$.
The inferred estimates of $N_{\Theta}$ are summarised in Table 7 for the fiducial cosmic shear analysis setup with all summary statistics, dubbed ?1cosmo?, as well as for the split cosmological analysis setups with COSEBIs considered in Sect. 4. For the split between angular scales, we employ band powers and 2PCFs, as discussed in Sect. 4.1.3. Additionally, we infer $N_{\Theta}$ for catalogue- and statistic-level splits in a single cosmological setup. For comparison, we provide the BMD, $N_{\Theta,{\rm BMD}}$, computed via Eq. (8). For both estimates, we list the resulting difference in $N_{\Theta}$ between the fiducial cosmological model and the split cosmological model, denoted as $d_{\chi^{2}}$ and $d_{\rm BMD}$. Overall, we found the number of constrained parameters to be higher in the split cosmological model. We attribute this to the doubling of the cosmological parameter space, where both instances of parameters sensitive to cosmic shear have been independently constrained.
Table 7: Effective number of constrained parameters. [^14] For each analysis setup we provide the effective number of constrained parameters inferred via $\chi^{2}$ minimisation of mock data vectors (Joachimi et al. 2021) in the third column, and the Bayesian model dimensionality (Handley & Lemos 2019a), inferred directly from the chain, in the fourth column. The first three rows show the results in an analysis with the fiducial setup, dubbed ‘1cosmo‘, in which we do not apply any splits to the data. The remaining rows present results when splitting the data into subsets. Here, we report the effective number of constrained parameters with the split cosmological model. For the splits at the catalogue level and the splits by summary statistic we additionally report the results with a shared set of parameters. The final two columns report the difference in the number of constrained parameters between the fiducial cosmological model and the split cosmological model. We employ COSEBIs as the summary statistic except when explicitly stated.
@@
Catalogue
North vs South 9.17 7.45 3.03 1.81
@@
Statistic
COSEBIs vs Band powers 10.22 10.69 3.26 4.59